OpenStudy (anonymous):

in a batch of 890 calculators, 12 were found to be defective. what is the probability that a calculator chosen at random will be percent. round to the nearest tenth of a percent if necessary.

OpenStudy (amistre64):

you might want to read your question again ....

OpenStudy (anonymous):

this one of my test question :(

OpenStudy (amistre64):

the question is meaningless as is ...

OpenStudy (anonymous):

a. 1.8% b. 1.3% c. 98.7% d. 74%

OpenStudy (anonymous):

defective=12/890=0.01348 ,0.0134*100=1.348 =1.3%